OpenStudy (anonymous):

Solve this equation. 4x + 5 = 21 2 4 12 16

OpenStudy (anonymous):

what grade are you in?

OpenStudy (saifoo.khan):

4x + 5 = 21 Bring 5 on the other side. 4x = 21 - 5 4x = 16 Divide 4 on both sides.

OpenStudy (anonymous):

That's got nothing to do with this question @zoejay1 and I think the answer is four @saifoo.khan